OpenStudy (oraclethinktank):

I need to graph the Quadratic Function y = -x^2 - 4x - 3

OpenStudy (oraclethinktank):

So far I have a= -x^2 b= -4x c= -3 Then, using x= -b/2a x= x^2/2(-4x) Which is equal to x^2/-8x. What comes next?

OpenStudy (oraclethinktank):

This is incorrect, For -b/2a, It should be 4x/2(-x^2) instead of what I originally put.

OpenStudy (oraclethinktank):

Doing the division, 4x/-2x^2= 2/x

OpenStudy (oraclethinktank):

Would that be the Line of Symmetry?

OpenStudy (oraclethinktank):

4x/-2x^2= - 2/x, rather.